OpenStudy (anonymous):

How do I rewrite cos^4 x so that it involves only the first power ?

OpenStudy (anonymous):

1-cos^4x then again use cos^2x-1=cos2x relation

OpenStudy (mathteacher1729):

One way of doing this... sin^2(x) = 1/2 (1-cos(2 x)) also anything^4 = (anything^2)^2 Hope this helps. :)